which is faster a bullet train with speed of 320 km/h or a race car speeding at 90 m/s?
1
Expert's answer
2020-08-27T10:25:09-0400

The speed of bullet train

"320\:\rm \frac{km}{hr}=320\times \rm \frac{1000\:m}{3600\: s}=88.9\: \frac{m}{s}"

Hence, the race car is faster than a bullet train.